OpenStudy (anonymous):

2^3-3*5-8=1 insert grouping symbols to make a true statement

OpenStudy (jack1):

\[2^3-3\times 5-8=-15\] (not 1) so \[(2^3 = 8) (-3\times5 = -15) (-8 = -8) =15 \] so try \[2^3-3\times(5-8)= 8 - (3\times-3)= 8-9 = -1\] hmm close but still not it maybe \[2^{(3-3)\times(5-8)}= 2^{(0)\times(-3)} = 2^0 = 1\] bingo!